a/b a=2/3 and b=5/6 I don't know how to do this. Can you help?
Answer by Alan3354(69443)

a/b a=2/3 and b=5/6
-------------
Invert the divisor and multiply
= (2/3)*6/5)
= 12/15
= 4/5